Minutes of the Trigger Operation Meeting (, 2007)
Attendees: Debbie Bard, Al Eisner, Kim Hojeong, Rahmat Rahmat, Su Dong.
- Quiet week - ROD is on-going and is expected to end Thursday.
- Emt object threshold study update here.
EMT object turn-on curves (plot 1a) show worrying sag between old and new data M curves - question of whether the objects are neutrals only or all EMT triggers? Binomial error bars would be useful, as would splitting the plot into barrel/endcap regions. The denominator may be different between old/new runs, could look at it's characteristics.
- L1TFmon tag in online release is old - Should newer tag be included in next release? Check with Jim which is used in LiveFastMon now.
- L1T now has correct cosmics reference plots in JAS. L3T will soon...
- The node disconnect problem mentioned last week happened again this week, and appears to have happened a couple of times a month since always. Not sure what causes this, but it's worrying!
Data Quality and Monitoring
- RQM is looking into the lumi plot problem now.
- Zk study - There are still problems getting this truth info out of L1TMooseApp, and problems getting the GLT info out of L1TKanApp. An impasse! Using L1TNtupleXtcApp and selecting L3 virtual compton events is also no good as the filter includes a cut of ~1GeV in energy, which is not useful for the 1Zk study where we want low momentum tracks. Instead, some event selection will have to be supplied, perhaps using two-phton selections.
This page is created by Debbie Bard (firstname.lastname@example.org); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.